A 60 Hz, three-phase. 6 – pole, Y-connected induction motor produces 100 hp at 1150 rpm. The

friction and windage losses of the motor are 1.2 kW, the stator copper and core losses are 2.1.kW

Compute the input power and the efficiency of the motor


1
Expert's answer
2021-08-05T03:19:53-0400

"Input \space power\\
P_{in}= \frac{T \omega}{63025}\\
P_{in}= \frac{2 \pi fN \omega}{63025}\\
P_{in}= \frac{2 \pi * 60*100 *1150}{63025}\\
P_{in}= 687.885\\"


"Efficiency \\
η= \frac{687.885-1.2-2.1}{687.885}*100\\
η= 99.5 \%"